The length of a rectangle is one unit more than its width. If the area is 30, square units find the dimensions of the rectangle.
Answer by checkley79(3341)

L=W+1
LW=AREA
(W+1)W=30
W^2+W=30
W^2+W-30=0
(W+6)(W-5)=0
W-5=0
W=5 ANS. FOR THE WIDTH.
L=5+1=6 ANS. FOR THE LENGTH.
PROOF:
6*5=30
30=30